# N-Agent Ad Hoc Teamwork

### v1

@@ 3 NAHT Problem Formulation
Let $g(k,\mu,s)$ represent a function for randomly selecting $k$ elements from a generic set $\mu$, conditioned on a state $s\in\mathcal{S}$. For instance, $g$ might be the distribution, Multinomial$(k,|\mu|,\bm{\phi}(s))$, for selecting $k$ elements with replacement from the set $\mu$, parameterized by the state-dependent probability logits $\bm{\phi}(s)\in[0,1]^{|\mu|}$. Allowing $g$ to depend on an initial state enriches the representable open interactions. Returning to our robot warehouse example, a state-dependent $g$ might enable us to model that robots suitable for heavy loads are more likely to be present near the loading dock area. For $s\in\mathcal{S}$, X is a sampling procedure parameterized by the tuple $(s,M,U,C,\bm{\phi}_{M},g_{U},g_{C})$, where $\bm{\phi}_{M}\in[0,1]^{M}$ represents the logits of a categorical distribution:
@@
Since teams consisting entirely of unknown agents are not of interest, we restrict consideration to teams containing at least one controlled agent, i.e $N\geq 1$. The agent sampling functions $g_{U},g_{C}$ could employ sampling with replacement to model scenarios where two robots may use the same policy, or sampling without replacement for scenarios where that is not possible (e.g. if all robots are heterogeneous).
Without loss of generality, let $C(\theta)$ denote a set of agents parameterized by $\theta$. Let the notation $\bm{\pi}^{(M)}\sim X(U,C)$ indicate sampling a team of size $M$ from $U$ and $C$. The objective of the NAHT problem is to find parameters $\theta$, such that $C(\theta)$ maximizes the expected return in the presence of teammates from $U$:
The central challenges of the NAHT problem include (1) coordinating with potentially unknown teammates (generalization), and (2) coping with an unknown number of uncontrolled teammates (openness).
4 A Motivating Example
Having introduced the NAHT problem, a natural question to consider is whether AHT solutions may optimally address NAHT problems as well. If so, then there would be little need to consider the NAHT problem setting. For instance, as an AHT solution consists of a single teammate policy, trained in the scenario illustrated in Figure 1, a simple yet reasonable approach consists of directly copying the AHT policy as many times as required, in an NAHT problem.
To illustrate the limitations of such an approach, we consider a concrete example in which a single ad hoc agent would follow a rational policy, but two ad hoc agents following that same policy would lead to a suboptimal outcome. Define the following simple game for $M$ agents: at each turn, each agent $a_{i}$ picks one bit $b_{i}\in\{0,1\}$; at the end of each turn, all the bits are summed $s=\sum_{i}b_{i}$. The team wins if and only if the sum of the chosen bits is exactly $1$.
Consider a scenario where there are $M-1$ (uncontrolled) agents who follow a deterministic policy. For simplicity, suppose they always pick 00. We now introduce a new ad hoc agent $a_{h}$. Assume that the ad hoc agent observes the $M-1$ always-zero agents play for one step, and the ad hoc agent $a_{h}$ is introduced into the game in the second step. Suppose the ad hoc agent selects actions according to the following strategy:
  1. 1.
    Model teammate actions by observing the last step.
  2. 2.
    Select the best response action.
  3. 3.
    If no information is known about a teammate, assume they act according to the average of expected actions of all the other observed agents.
If the ad hoc agent $a_{h}$ acts based on the teammate modeling scheme described above, it will observe that all the other agents picked 00 at the first step, and therefore pick $1$.
Now, consider a scenario in which there are $M-2$ always-zero agents playing without the ad hoc agent for one step, and the ad hoc agents $a_{h}^{1}$ and $a_{h}^{2}$ are introduced into the game in the second step. Following the same modeling policy, the two ad hoc agents $a_{h}^{1}$ and $a_{h}^{2}$ will both assume that their teammates follow an always-zero policy, and therefore learn to pick the 1 action—a suboptimal decision, since $s=2$. Following the second step, both $a_{h}^{1}$ and $a_{h}^{2}$ will assume that their counterparts pick the $1$ action. As a result, both agents will switch to the action they would consider the best response, and would both pick the 00 action, meaning once again the joint action is not optimal, as $s=0$. We observe that extending the agent modeling window beyond one step does not resolve this inherent pattern (see Appendix A.4).
While the policy above leads to optimal behavior for a single ad hoc agent, it results in poor joint performance in the presence of another ad hoc agent. A policy that explicitly accounts for the potentially unpredictable behavior of other agents and strives to cooperate with other agents would lead to better performance. In the example above, finding such a policy is as simple as introducing $\epsilon$-greedy exploration to the ad hoc agents’ action selection behavior. With that modification, both agents will continue exploring until the optimal reward of $s=1$ is achieved, in which case both agents will converge to an optimal behavior. This is because with high probability, the two ad hoc agents will learn to optimally coordinate by following a joint policy that selects one of the optimal joint actions of $(0,1)$ or $(1,0)$ at some time step, after which both agents will repeat that joint action due to the single-step agent modeling.
Figure 2: The training protocol for the proposed method, POAM. POAM adopts an independent policy gradient architecture with parameter sharing. POAM trains a single policy network, which characterizes the behavior of all controlled agents. The yellow agents are controlled by policies drawn from $U$, the set of uncontrolled agents, while the green agents are controlled by the same POAM policy. Data from both controlled and uncontrolled agents is used to train the POAM critic, $V_{\pi}$, while the actor $\pi(\theta)$ is trained on data from the controlled agents only. The actor and critic are both conditioned on a learned team encoding vector, learned via an encoder-decoder architecture.
5 Policy Optimization with Agent Modeling (POAM)
This section describes the proposed Policy Optimization with Agent Modeling (POAM) method, which trains a collection of NAHT agents that can adaptively deal with different collections of unknown teammates. POAM relies on an agent modeling network to initially build an encoding vector of the collection of teammates encountered during an interaction. Adaptive agent policies that can maximize the controlled agents’ returns are then learned by POAM by training a policy network, which receives environment observation and team encoding vector as input. Details of the training process for both agent modeling and policy networks are described in Sections 5.1 and 5.2 respectively, while an illustration of how POAM trains NAHT agents is provided in Figure 2.
5.1 Agent Modeling Network
Designing adaptive policies that enable NAHT agents to achieve optimal returns against any team of uncontrolled agents drawn from some set $U$, requires information on the encountered team’s unknown behavior. However, in the absence of prior knowledge about uncontrolled teammates’ policies, local observations from a single timestep may not contain sufficient information regarding the encountered team. To circumvent this lack of information, POAM’s agent modeling network plays a crucial role in providing team encoding vectors that characterize the behavior of teammates in the encountered team.
We identify two main criteria for desirable team encoding vectors. First, team encoding vectors should identify information regarding the unknown state and behavior of other agents in the environment. Second, team encoding vectors should ideally be computable from the sequence of local observations and actions of the team. Fulfilling both requirements provides an agent with useful information for decision-making in NAHT problems, even under partial observability.
POAM produces desirable team encoding vectors by training a model with an encoder-decoder architecture, which is illustrated by red components in Figure 2. The encoder, $f^{\text{enc}}_{\theta^{e}}:H^{i}\mapsto\mathbb{R}^{n}$, is parameterized by $\theta^{e}$ and processes the modeling agent’s history of local observations and actions up to timestep $t$, $h^{t}=\{o_{i}^{k},a_{i}^{k-1}\}_{k=1}^{t}$, to compute a team encoding vector, $e^{t}_{i}\in\mathbb{R}^{n}$. This reliance on local observations helps ensure that the agent modeling network can operate without having access to the environment state (i.e. under partial observability). We then train decoder networks, $f^{\text{dec}}_{\theta^{o}}:\mathbb{R}^{n}\mapsto O^{-i}$ and $f^{\text{dec}}_{\theta^{a}}:\mathbb{R}^{n}\mapsto\Delta(A_{-i})$, that predicts the observations and actions of the modelled agents at timestep $t$, $(o_{-i}^{t},a_{-i}^{t})$, to encourage $e^{t}_{i}$ to contain relevant information for the NAHT agents’ decision-making process. Formally, this decoder is trained to minimize the following loss function to accurately predict the modelled agents’ observations and actions:
5.2 Policy and Value Networks
POAM relies on an actor-critic architecture to train agent policies, where the policy and networks are both conditioned on the teammate embedding produced by the encoder-decoder architecture described in Section 5.1.
POAM then trains the policy and value networks using an approach based off the Independent PPO algorithm (Yu et al., 2022) (IPPO). IPPO is selected as the base MARL algorithm for two reasons. First, using an independent MARL method circumvents the need to deal with the changing number of agents resulting from environment openness. Second, it has been demonstrated to be effective on various MARL tasks. POAM trains the value network to produce accurate state value estimates by minimizing the following loss function:
where $\hat{V}^{t}=\sum_{l=0}^{T}\gamma^{l}r^{t+l}$. The policy network is trained to minimize the PPO loss function, where the Generalized Advantage Estimator is employed (Schulman et al., 2015):
$L_{\theta^{p}}(h^{t})=\text{min}\left(r^{t}(\theta^{p})\hat{A}^{t}_{GAE},\text{clip}\left(r^{t}(\theta^{p}),1-\epsilon,1+\epsilon\right)\hat{A}^{t}_{GAE}\right)-\alpha{H}(\pi^{i}_{\theta^{p}}(.|h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))),$ (4)
where $r^{t}(\theta^{p})=\dfrac{\pi^{i}_{\theta^{p}}(a^{t}_{i}|h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))}{\pi^{i}_{\theta^{p_{old}}}(a^{t}_{i}|h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))}$.
Leveraging data from uncontrolled agents
In the NAHT setting, we assume access to the joint observations and actions generated by the current team deployed in the environment, where the team consists of a mix of controlled and uncontrolled agents. POAM leverages the data from both controlled and uncontrolled agents to train the value network—in effect, treating the uncontrolled agents as exploration policies. This is a significant departure from PPO, which is a fully on-policy algorithm. However, only data from the controlled agents is used to train the policy network, as the policy update is highly sensitive to off-policy data.
6 Experiments and Results
This section presents an empirical evaluation of POAM and baseline approaches across different NAHT problems. We specifically investigate two questions. First, whether POAM leads to improved solutions for NAHT over baselines, in terms of sample efficiency or asymptoptic task return. Second, whether agent modeling enables generalization to previously unseen teammates.
We start this section by introducing the main experimental procedures and baselines in Section 6.1. Section 6.2 presents our main result that compares POAM and the baselines in NAHT problems situated across different environments. Sections 6.36.4 then provide additional analysis on the team encoding vectors produced by POAM and its ability to generalize the learned policy to previously unseen sets of teammate policies. Finally, see the Appendix for further results, hyperparameters, and implementation details. Our code will be made available upon publication.
6.1 Experimental Design
This subsection describes the experimental design, including the particular NAHT instance considered in the experiments, training procedure, experimental domain, and baselines.
6.1.1 A Practical Instantiation of NAHT
@@
Similar to AHT, the NAHT problem can be parameterized by the amount of knowledge that controlled agents have about uncontrolled agents, and whether uncontrolled agents can adapt to the behavior of controlled agents (Barrett & Stone, 2012). This section considers a practical instantiation of the NAHT problem, designed to assess whether the baseline algorithms have successfully learned to cooperate with each of the uncontrolled agents under consideration.
@@
Recall that $U$ and $C$ denote the sets of uncontrolled and controlled agent policies, respectively, while $M$ denotes the total team size, and $N$ the number of agents sampled from $C$. For simplicity, we assume that policies in $U$ do not adapt to the controlled agents’ behavior, agents are homogeneous (so that they are interchangeable within a team), and that agents cannot communicate. While in principle, the team sampling procedure $X$ could define any probability distribution over teams constructed from $U$ and $C$, we consider an $X_{train}$ consisting of sampling $N$ uniformly from $\{1,\cdots,M-1\}$, sampling $N$ agents from $C$ and $M-N$ agents from $U$ in a uniform fashion (Figure 3). The sampling procedure described here takes place at the beginning of each episode to select a team, which is then deployed in the environment. Data generated by the deployed team (e.g. joint observations, joint actions, and rewards) are returned to the learning algorithm. See Appendix A.1.2 for further details on the evaluation procedure.
6.1.2 Generating Uncontrolled Teammates
To generate a set of uncontrolled teammates $U$, the following MARL algorithms are used to train agent teams: VDN (Sunehag et al., 2018), QMIX (Rashid et al., 2018), IQL (Tan, 1997), IPPO, and MAPPO (Yu et al., 2022). We verify that the generated team behaviors are diverse by checking that (1) teams trained by the same algorithm learn non-compatible coordination conventions, and (2) teams trained by different algorithms are not compatible (see Appendix A.2).
6.1.3 Experimental Domain
Experiments are conducted on the StarCraft Multi-Agent Challenge (SMAC) benchmark (Samvelyan et al., 2019). SMAC features a set of cooperative tasks, where a team of allied agents must defeat a team of enemy agents controlled by the game server. It is a partially observable domain with a continuous state space, where each agent can observe features about itself, enemy, and allied agents within some radius. The action space is discrete—each agent can choose an enemy to attack, a direction to move in, or to not perform any action. The SMAC tasks considered here are 5v6, 8v9, 10v11, 3s4z, 3s5z. The first number indicates the number of allied agents, while the second indicates the number of enemy agents; full task names and descriptions are provided in App. A.1.
6.1.4 Baselines
As NAHT is a new problem proposed by this paper, there are no prior algorithms that are directly designed for the NAHT problem. Therefore, we construct three baselines to contextualize the performance of POAM.
  • Naive MARL: various well-known MARL algorithms are considered, including both independent and centralized training with decentralized execution algorithms (Hernandez-Leal et al., 2019). The algorithms evaluated here include IQL (Claus & Boutilier, 1998), VDN (Sunehag et al., 2018), QMIX Rashid et al. (2018), IPPO, and MAPPO (Yu et al., 2022). The MARL baselines are trained in self-play and then evaluated in the NAHT setting. In the following, we report only the performance of the best naive MARL baseline.
  • Independent PPO in the NAHT setting (IPPO-NAHT): IPPO is a policy gradient MARL algorithm that directly generalizes PPO (Schulman et al., 2017) to the multi-agent setting. It was found to be surpringly effective on a variety of MARL benchmarks by Yu et al. (2022). In contrast to the naive MARL baselines, IPPO-NAHT is trained using the proposed NAHT training scheme. The variant considered here employs full parameter sharing, where the actor is trained on data only from controlled agents, but the critic is trained using data from both controlled and uncontrolled agents. The latter detail is a key algorithmic feature which POAM also employs, but is an extension from the most naive version of PPO (see Section 6.3). IPPO can be considered an ablation of POAM, where the agent modeling module is removed.
  • POAM in the AHT setting (POAM-AHT): As considered in Section 4, a natural baseline approach to the NAHT problem is to deploy AHT policies directly in the NAHT setting. To experimentally evaluate the intuition that AHT policies do not suffice for the NAHT problem setting, we consider an AHT version of POAM. POAM-AHT is trained identically to POAM, but where the number of controlled agents $N$ is always equal to 1 during training.
All methods employ full parameter sharing (Papoudakis et al., 2021b). See Appendix A.1 for details about architecture, hyperparameter tuning and algorithm implementation.
6.2 Main Results
Figure 4: POAM consistently improves over the baselines of IPPO-NAHT, POAM-AHT, and the best naive MARL baseline in all tasks, in either sample efficiency or asymptotic return. The performance of POAM-AHT and naive MARL are inconsistent across tasks. Results are computed over three trials, while the shaded region represents the 95% confidence interval.
The primary question addressed here is whether POAM improves performance on the NAHT problem setting, as compared to baselines. Figure 4 shows the learning curves of POAM and IPPO-NAHT on the SMAC tasks, and the test returns achieved by the best naive MARL baseline and POAM-AHT. See the Appendix for tables providing the performances of all naive MARL methods, across all tasks. All learning curves consist of the mean test returns across three trials, while the shaded regions reflect the 95% confidence intervals. All methods are trained for 20M time steps. We find that for all tasks, POAM outperforms baselines in either asymptotic return (5v6, 3s5z), or sample efficiency (8v9, 10v11). While POAM initially has lower sample efficiency than IPPO-NAHT, this trend is reversed after the first few million steps, after which POAM displays higher return. We attribute this to the initial costs incurred by learning team encoding vectors, which once learned, improves learning efficiency. IPPO-NAHT, which can be viewed as an ablation of POAM with no agent modeling, is the next best performing method. Although IPPO-NAHT has generally poorer sample efficiency than POAM, the method converges to the approximately the same returns on 8v9 and 10v11. While the agent modeling module of POAM provides team embedding vectors to the policy learning process, the embeddings are themselves produced from each agent’s own observation. Since no additional information is provided to POAM agents, it is unsurprising that IPPO-NAHT can converge to similar solutions as POAM, given enough training steps.
Figure 5: Evolution of the within-episode mean squared error (left) and probability of actions that were actually taken by modelled teammates (right), computed by the POAM agent over the course of training. The agent modeling performance improves over the course of an episode, as more data about teammate behavior is observed.
Finally, while the best naive MARL baseline and POAM-AHT learn good solutions on some tasks, neither method consistently performs well across all tasks. Overall, POAM discovers the most consistently performant policies compared to the other methods. We conclude that (1) the proposed learning scheme can learn effectively in the presence of multiple types of agents, and (2) agent modeling improves learning efficiency.
6.3 A Closer Look at POAM
Two key aspects of POAM are the use of a recurrent teammate modeling module, and the use of data from uncontrolled agents to train the critic. We perform further analysis of both of these aspects on the 5v6 task.
Teammate modeling performance
A natural question then, is whether the learned teammate encoding vectors actually improve over the course of an episode. Figure 5 depicts the within-episode mean squared error (MSE) of the observation predicted by the decoder, and the within-episode probability of the action actually taken by the modelled teammates, according to the decoder’s modelled action distribution.
Each curve corresponds to a different training checkpoint. We observe the MSE rapidly converges to a low value, while the predicted probability of the taken actions has converged to 1 by time step 20. Note that the MSE is initially 0 at timestep 0 because the initial position is not randomized in SMAC. Next, we observe that the modeling performance improves over the course of training, indicated by lower prediction error at later training steps. Thus, we conclude that POAM learns accurate teammate models relatively early in the episode for the 5v6 task—successfully coping with the challenges introduced by the sampled teammates and partial observability.
Impact of data from non-controlled agents
Recall that both POAM and IPPO-NAHT update the value network using data from both the controlled and uncontrolled agents. As Figure 6 shows, this algorithmic feature results in a significant performance gain over training using on-policy data only, for both POAM and IPPO-NAHT.
6.4 Out of Distribution Generalization
Experiments in prior sections were conducted with in-distribution teammates, i.e. teammates that the methods interacted directly with during training. This section examines whether agent modeling improves generalization to out-of-distribution (OOD) teammates, where the OOD teammates are generated by using MARL algorithms with different random seeds to generate teammates that have not been interacted with during training. [^4] We demonstrate in App. A.2 that in SMAC tasks, the same MARL algorithm with different seeds trains teams that have different coordination conventions. Hence, teams from unseen seeds display OOD behavior.
Figure 7 shows the mean and standard error of the test return achieved by POAM, compared to IPPO-NAHT, when the algorithm in question is paired with previously unseen seeds of IPPO, IQL, MAPPO, QMIX, and VDN (computed across three trials). In two out of four tasks (5v6, 3s5z), POAM has a significantly higher return than IPPO-NAHT, while the remaining two tasks exhibit a smaller improvement. For all tasks, we observe that the standard error for POAM is lower than that of IPPO-NAHT, showing that POAM performs more consistently against a range of teammates and team configurations than IPPO-NAHT.
7 Related Works
Ad Hoc Teamwork & Zero-Shot Coordination.
Prior works in ad hoc teamwork (Stone et al., 2010) and zero-shot coordination (ZSC) (Hu et al., 2020) explored methods to design adaptive agents that can optimally collaborate with unknown teammates. While they both highly resemble the NAHT problem, existing methods for AHT (Rahman et al., 2021; Mirsky et al., 2022) and ZSC (Hu et al., 2020; Lupu et al., 2021) have been limited to single-agent control scenarios. We argue that direct, naive applications of AHT and ZSC techniques to our problem of interest are ineffective—see the discussion in Section 4 and results in Section 6.
Recent research in AHT and ZSC utilizes neural networks to improve agent collaboration within various team configurations. These recent works mostly focus on two approaches. The first approach trains the agent to adapt to unknown teammates by characterizing teammates’ behavior as fixed-length vectors using neural networks and learning a policy network conditioned on these vectors (Rahman et al., 2021; Papoudakis et al., 2021a; Zintgraf et al., 2021). The second designs teammate policies that maximize the agent’s performance when collaborating with diverse teammates (Lupu et al., 2021; Rahman et al., 2021). Our work builds on the first category, extending it to control multiple agents. While this also offers a potential path for robust NAHT agents, designing teammate policies for training will be kept as future work.
Evaluating Agents’ Cooperative Capabilities.
Beyond training agents to collaborate with teammates having unknown policies, researchers have developed environments and metrics to assess cooperative abilities. The Melting Pot suite (Leibo et al., 2021; Agapiou et al., 2022) mostly evaluates controlled agents’ ability to maximize utilitarian welfare against unknown agents. However, this evaluation suite focuses on mixed-motive games where agents may have conflicting goals. This contrasts with our work’s scope of fully cooperative settings, where all agents share the same reward function.  MacAlpine & Stone (2017) also explored alternative metrics to measure agents’ cooperative capabilities while disentangling the effects of their overall skills in drop-in RoboSoccer.
Agent Modeling.
Agent modeling (Albrecht & Stone, 2018) helps us characterize other agents based on their actions. Such characterizations could attempt to infer important attributes of another agent’s decision making process such as their goals (Hanna et al., 2021) or policies (Papoudakis et al., 2021a). POAM and most existing AHT techniques rely on agent modeling to provide important teammate information for decision-making when collaborating with unknown teammates.
Cooperative MARL.
Cooperative Multi-Agent Reinforcement Learning (MARL) explores algorithms for training agent teams on fully cooperative tasks. Some existing methods focus on credit assignment and decentralized control (Foerster et al., 2018; Rashid et al., 2018). Other works in cooperative MARL also leverage parameter sharing and role assignment (e.g., Christianos et al. (2021); Yang et al. (2022)) to decide an optimal division of labor between agents. However, these techniques assume control over all existing agents during training and evaluation. This limits these techniques’ effectiveness in settings with new teammates (NAHT problem) as shown in prior work (Vezhnevets et al., 2020; Hu et al., 2020; Rahman et al., 2021).
8 Conclusion
This paper proposes and formulates the problem of $N$-agent ad-hoc teamwork (NAHT), a generalization of both AHT and MARL. It further proposes a multi-agent reinforcement learning algorithm to train NAHT agents called POAM, and develops a procedure to train and evaluate NAHT agents. POAM is a policy gradient approach that uses an encoder-decoder architecture to perform teammate modeling, and leverages data from uncontrolled agents for policy optimization. Empirical validation on StarCraft II tasks shows that POAM consistently improves over baseline NAHT approaches, in terms of sample efficiency, asymptotic return, and generalization to out-of-distribution teammates.
While this work explores teammate modeling for NAHT, other algorithmic ideas from AHT provide a rich set of future directions. These include generating teammates for the NAHT setting, alternative teammate modeling schemes, and considering communication protocols for NAHT agents. Limitations of POAM also suggest directions for future work. For instance, POAM is an independent MARL method (albeit with parameter sharing)—how might one leverage knowledge from the global state to generate improved NAHT policies? Further, POAM’s actor update is purely on-policy, and therefore cannot leverage data generated by uncontrolled agents. Future work may consider off-policy learning methods as a solution strategy.
Acknowledgments
This work has taken place in the Learning Agents Research Group (LARG) at UT Austin. LARG research is supported in part by NSF (FAIN-2019844, NRT-2125858), ONR (N00014-18-2243), ARO (E2061621), Bosch, Lockheed Martin, and UT Austin’s Good Systems grand challenge. Peter Stone serves as the Executive Director of Sony AI America and receives financial compensation for this work. The terms of this arrangement have been reviewed and approved by the University of Texas at Austin in accordance with its policy on objectivity in research.
Appendix A Appendix
A.1 Experiment Details
A.1.1 SMAC Tasks

### v2

@@ A.1 An Example of the Team Sampling Procedure
Let $g(k,\mu,s)$ represent a function for randomly selecting $k$ elements from a generic set $\mu$, conditioned on a state $s\in\mathcal{S}$. For instance, $g$ might be the distribution, Multinomial$(k,|\mu|,\bm{\phi}(s))$, for selecting $k$ elements with replacement from the set $\mu$, parameterized by the state-dependent probability logits $\bm{\phi}(s)\in[0,1]^{|\mu|}$. Allowing $g$ to depend on an initial state enriches the representable open interactions. Returning to our robot warehouse example, a state-dependent $g$ might enable us to model that robots suitable for heavy loads are more likely to be present near the loading dock area. Similarly, if the time is included in the state, we would be able to model dynamic characteristics, e.g. that certain types of robots are only available during regular working hours, when humans are available to supervise. Of course, in simple cases, $g$ might be state-independent. For $s\in\mathcal{S}$, $X$ is a sampling procedure parameterized by the tuple $(s,M,U,C,\bm{\phi}_{M},g_{U},g_{C})$, where $\bm{\phi}_{M}\in[0,1]^{M}$ represents the logits of a categorical distribution:
@@
Note that the agent sampling functions $g_{U},g_{C}$ could employ sampling with replacement to model scenarios where two robots may use the same policy, or sampling without replacement for scenarios where that is not possible (e.g. if all robots are heterogeneous).
A.2 Further Discussion of the Motivating Example
This section provides proofs for claims in Section 4, and further discussion in support of the motivating example.
Lemma A.1.
For a team of $M$ agents independently and identically selecting actions with probability $p$, the $p$ that maximizes the probability of winning is $p=\frac{1}{M}$.
Proof.
The probability of winning, $P(s=1)$ may be computed as follows:
$P(s=1)=M*P(\text{agent }i\text{ chooses 1 and $i^{-}$ choose 0})=Mp(1-p)^{M-1}.$
We wish to analytically compute the global maxima of this expression with respect to $p$, by considering the boundary points $p\in{0,1}$ and the zeros of the first derivative.
Clearly, if all agents select 1 with probability 0, the team cannot win; thus, $p=0$ is not a maxima. Similarly, if all agents select 1 with probability 1, the team also never wins; thus, $p=1$ is also not a maxima. Next, we compute the derivative of $P(s=1)$:
$\frac{d[P(s=1)]}{dp}$ $=M[(1-p)^{M-1}-(M-1)p(1-p)^{M-2}]$
$=M[(1-p)^{M-2}[(1-p)-(M-1)p]].$
The zeros of the above expression occur at $p=0$ and $p=\frac{1}{M}$. For $p=\frac{1}{M}$, note that $P(s=1)=(\frac{M-1}{M})^{M-1}>0$; thus, $p=\frac{1}{M}$ must be the global maximum. ∎
Lemma A.2.
In a team of three agents consisting of two uncontrolled agents who select 1 with probability $p=\frac{1}{3}$, and one controlled ad hoc agent who selects 1 with probability $p_{\text{aht}}$, the probability of winning is $P(s=1)=\frac{4}{9}$.
Proof.
Let $s_{u}$ denote the sum of the bits chosen by the two uncontrolled agents, and $b_{aht}$ denote the bit value chosen by the controlled agent. The probability of winning can be computed by partitioning the winning outcomes by whether $s_{u}=1$ and the ad hoc agent selects 0, or whether the ad hoc agent selects 1 and both uncontrolled agents select 0, $s_{u}=0$.
$P(s=1)$ $=P(s_{u}=1\land b_{aht}=0)+P(s_{u}=0\land b_{aht}=1)$
$=2\cdot\frac{1}{3}\cdot\frac{2}{3}\cdot\bigl{(}1-p_{\text{aht}}\bigr{)}+p_{\text{aht}}\left(\frac{2}{3}\cdot\frac{2}{3}\right)$
$=\frac{4}{9}-\frac{4}{9}\cdot p_{\text{aht}}+\frac{4}{9}\cdot p_{\text{aht}}$
$=\frac{4}{9}.$
Lemma A.3.
In a team of three agents where two ad hoc agents select 1 with probability $p_{\text{aht}}$ and the remaining uncontrolled agent selects 1 with probability $\frac{1}{3}$, the maximizing $p_{\text{aht}}=\frac{1}{3}$ and the corresponding winning probability is $P(s=1)=\frac{4}{9}$.
Proof.
Let $s_{aht}$ denote the sum of the bits chosen by the two ad hoc agents, and $b_{u}$ denote the bit value chosen by the uncontrolled agent. The probability of winning may be directly computed as follows:
$P(s=1)$ $=P(s_{aht}=1\land a_{u}=0)+P(s_{aht}=0\land a_{u}=1)$
$=2\cdot p_{\text{aht}}\cdot(1-p_{\text{aht}})\cdot\frac{2}{3}+(1-p_{\text{aht}})^{2}\cdot\frac{1}{3}$
$=(1-p_{\text{aht}})\left[\frac{4}{3}p_{\text{aht}}+\frac{1}{3}(1-p_{\text{aht}})\right]$
$=(1-p_{\text{aht}})\left(\frac{1}{3}+p_{\text{aht}}\right).$
To determine the maximizing value, we compute $P(s=1)$ at the boundary points $p_{\text{aht}}\in\{0,1\}$ and at the zeros of the derivative of $P(s=1)$. Note that for $p_{\text{aht}}=0$, $P(s=1)=\frac{1}{3}$, while for $p_{\text{aht}}=1$, $P(s=1)=0$.
The derivative of the analytic expression of $P(s=1)$ with respect to $p_{\text{aht}}$ is:
$\frac{d[P(s=1)]}{dp_{\text{aht}}}=(1-p_{\text{aht}})-\left(\frac{1}{3}+p_{\text{aht}}\right)=\frac{2}{3}-2p_{\text{aht}}.$ (4)
The zeros of the above expression occur at $p_{\text{aht}}=\frac{1}{3}$, which corresponds to $P(s=1)=\frac{4}{9}$. ∎
Lemma A.4.
In a team of three agents where one agent always selects 0, one agent selects 1 with probability $\frac{1}{3}$, and the last agent selects 1 with probability $p_{\text{naht}}$, the optimal $p_{\text{naht}}=1$ and results in a winning probability of $P(s=1)=\frac{2}{3}$.
Proof.
Since one of the two controlled agents always plays 0, the game effectively becomes a two-player game instead.
Let $a_{u}$ denote the action selected by the uncontrolled agent, and let $a_{naht}$ denote the action of the controlled agent. The probability of winning may be directly computed as follows:
$P(s=1)$ $=P(a_{u}=1\land a_{naht}=0)+P(a_{u}=0\land a_{naht}=1)$
$=\frac{1}{3}\cdot(1-p_{\text{naht}})+\frac{2}{3}\cdot p_{\text{naht}}$
$=\frac{1}{3}+\frac{1}{3}\cdot p_{\text{naht}}.$
It is clear that $p_{\text{naht}}=1$ maximizes $P(s=1)$, and the corresponding value of $P(s=1)=\frac{2}{3}$. ∎
A.3 Experiment Details
A.3.1 Team Sampling Procedure Used in Experiments
@@
Recall that $U$ and $C$ denote the sets of uncontrolled and controlled agent policies, respectively, while $M$ denotes the total team size, and $N$ the number of agents sampled from $C$. The experiments in the following consider an $X_{train}$ consisting of sampling $N$ uniformly from $\{1,\cdots,M-1\}$, sampling $N$ agents from $C$ and $M-N$ agents from $U$ in a uniform fashion (Figure 7). The sampling procedure takes place at the beginning of each episode to select a team, which is then deployed in the environment. Data generated by the deployed team (e.g. joint observations, joint actions, and rewards) are returned to the learning algorithm. See Appendix A.3.3 for further details on the evaluation procedure.
A.3.2 Experimental Domains
MPE Predator Prey
The MPE environment [23] (released under the MIT License) is a setting where particle agents interact within a bounded 2d plane, equipped with a discrete action space and continuous observation space. The observation space is 16-dimensional, and contains the agents’ own position and velocity, relative positions/velocities of all other agents, landmarks, and prey. The action space consists of 5 discrete actions, corresponding to the four cardinal movement directions and a no-op action.
The predator-prey task (mpe-pp) is a custom task implemented by the authors of this paper within the fork of the MPE environment released by Papoudakis et al. [29] (MIT License). In this task, three predators must cooperatively pursue a pre-trained prey agent. The team receives a reward of 1 if two or more predators collide with the prey at a single time step, and no reward if only a single agent collides with the prey. A shaping reward consisting of 0.01 times the $\ell_{2}$ distance between each agents in the team and the prey is provided as well. The maximum episode length is 100 time steps.
SMAC
SMAC [33] (released under the MIT License) features a set of cooperative tasks, where a team of allied agents must defeat a team of enemy agents controlled by the game server. It is a partially observable domain with a continuous state space and discrete action space. For each agent, the observation space is continuous, and consists of features about itself, enemy, and allied agents within some radius. The action space is discrete, and allows an agent to choose an enemy to attack, a direction to move in, or to not perform any action. As the number of allies and enemies varies between tasks, the dimensionality of the observation and action spaces is particular to each task. The maximum number of time steps per episode is also task specific, although early termination occurs if all enemies are defeated. At each time step, the team receives a shaped reward corresponding to the damage dealt, and bonuses of 10 and 200 points for killing an enemy and winning the scenario by defeating all enemies. The reward is scaled such that the maximal return achievable in each task is 20.